OpenStudy (anonymous):

complete the ordered triple 8x-y+z=32 3x+2y-3z=36 x-3y+2z=-14

OpenStudy (amistre64):

piece o cake... :)

OpenStudy (amistre64):

how you wanna solve it?

OpenStudy (anonymous):

i thought so too but when i worked the equation the way i was explained didnt get the right answer

OpenStudy (anonymous):

directions say to number equations 1-3 multiply (1) by 2

OpenStudy (amistre64):

8x - y + z =32(-3/8) 8x - y + z =32(-1/8) 3x +2y -3z =36 x -3y +2z =-14

OpenStudy (anonymous):

what is that? confusing me more

OpenStudy (amistre64):

We are multiplying row 1 by a number that will turn the first term into a negative of the equation underneath it. -3x +3/8 y -3/8z = -96/8 -1x +1/8 y -1/8z = -32/8 3x + 2y -3 z = 36 x -3 y +2 z =-14 --------------------- --------------------

OpenStudy (amistre64):

-3x + 3/8 y -3/8z = -12 -1x +1/8 y - 1/8z = - 4 3x +16/8y -24/8z = 36 x -24/8 y +16/8z =-14 --------------------- ---------------------- 0 +19/8y -27/8z = 24 0 -23/8y +15/8z = -28 and we end up with two equations of two variables that are easier to work with....

OpenStudy (anonymous):

the way professor taught me is multiply equation (1) by 2 16x-2y+2z=64 +3x=2y-3z=36 equation (2) =19x-z=100 which becomes equation (4)

OpenStudy (amistre64):

your professor choose to eliminate the "y" variable; I simply choose to start at the front and aliminatte the "x" variable. Same rules, different variable.

OpenStudy (anonymous):

ok his way seems less confusing

OpenStudy (amistre64):

his way keeps the fractions out :) but the process is the same. we use the "modified" equation to eliminate a variable from the other 2

OpenStudy (anonymous):

oh good i hate fractions

OpenStudy (amistre64):

once you know the process..... the way you get to it is up to you :)

OpenStudy (amistre64):

there is no special rule about : you have to use equation 1 and multiply by a certain variable yadayadayada.... pick your poison and drink ;)

OpenStudy (anonymous):

good way to put it

OpenStudy (anonymous):

can we work on a different one

OpenStudy (amistre64):

we can :)

OpenStudy (anonymous):

x+y+z=4 2x+4y+2z=4 -x+8y-3z=-34 please leave out the fractions on my end anyway

OpenStudy (amistre64):

:) lets use equation 1 since it looks the plainest and use it to "calibrate" the other 2. ok?

OpenStudy (amistre64):

x + y + z =4 <- what should we mulitply this bu to get rid 2x+4y+2z =4 of a variable in the bottom equation? -------------

OpenStudy (amistre64):

we want either a: -2x, or a -4y; or even a -2z... what variable you wanna destroy?

OpenStudy (amistre64):

x y or z.... just pick one so we can move ahead ;)

OpenStudy (anonymous):

x

OpenStudy (amistre64):

good, we will destroy the "x" :)

OpenStudy (amistre64):

the top has an "x" and the bottom has an "2x. ___ x + 2x = 0..... what do we need to fill in the blank here?

OpenStudy (anonymous):

your doing it the hard way

OpenStudy (amistre64):

lol..... no I aint :)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

sry 2

OpenStudy (amistre64):

so you want me to believe that: 1x + 2x = 0?

OpenStudy (amistre64):

2x + 2x = 0?

OpenStudy (anonymous):

-2

OpenStudy (amistre64):

yay!!.... we hav a winner ;) -2 it is.

OpenStudy (amistre64):

x + y + z =4 <- multiply everything in this top by (-2) 2x+4y+2z =4 to get -2x -2y -2z =-8 2x +4y +2z =4 ; now add them together. ----------------- 0 +2y +0 = -4 ; wow, 2 variables with one shot lol

OpenStudy (anonymous):

question? y do we put - between 2y, and 2z

OpenStudy (anonymous):

because we multiply by -2, tina duh

OpenStudy (amistre64):

i dont understand the question... yeah, thats the ticket :)

OpenStudy (amistre64):

so we get; 2y = -4; what does y equal since its all alone?

OpenStudy (anonymous):

better question why do we multiply by -2

OpenStudy (amistre64):

becasue you decided to destroy the x variables...... and that meant that we needed to have a number such that: ___x + 2x = 0 ; zero means they are eliminated, destroyed, kaput right?

OpenStudy (anonymous):

ok got it

OpenStudy (amistre64):

(__)x + 2x = 0 ; you recall that :)

OpenStudy (amistre64):

you sophies choiced it lol

OpenStudy (anonymous):

going for bs in accounting didnt know this kind of math would be part of it

OpenStudy (anonymous):

im actually a freshman

OpenStudy (amistre64):

gotta know algebra... and thats all this is :)

OpenStudy (anonymous):

still trying to figure out how i made it out of hs

OpenStudy (amistre64):

So, to move ahead... we came to the conclusion that both those variables vanished in the process, and we are left with 2y=-4; whats y have to be?

OpenStudy (anonymous):

-2

OpenStudy (amistre64):

good, write that down someplace...y = -2 its one of your answers :)

OpenStudy (amistre64):

Now we use equation (1) that you gave me and use it to eliminate "x" in the last one right? x + y + z = 4 -x +8y -3z =-34

OpenStudy (amistre64):

Is there anything we have to do to this to get: x - x = 0?

OpenStudy (anonymous):

do we multiply equation 1by -3?

OpenStudy (anonymous):

1-1=0

OpenStudy (amistre64):

no..... you chose "x" variable for elimination...stick with it to the end :)

OpenStudy (anonymous):

ok, carry on then

OpenStudy (amistre64):

(__)x -x = 0 ; what do we have to do to make this right? what number do we need to use?

OpenStudy (amistre64):

that format lied to me lol

OpenStudy (anonymous):

? idk x represents 1 right?

OpenStudy (amistre64):

x represents x.... 1 represents 1..... but what you a really asking is: isnt there an invisible 1 in front of those? and I say , yes. and so you say, then the number we need is 1; and I say correct :)

OpenStudy (anonymous):

yeah thats exactly what i was thinking

OpenStudy (amistre64):

then we can multiply everything in the top by (1)...which seems pretty pointless right? but thats what we do :) x + y + z =4 <-- multiply it all by (1) and we get. -x +8y -3z =-34 x+ y+ z =4 -x+8y-3z=-34 <-- add em up --------------- 0 +7y -2z = -30

OpenStudy (amistre64):

ummm 8 + 1 = 9 :) I forgot how to add :)

OpenStudy (amistre64):

8y -2z = -30 now we can use this equation for some good :)

OpenStudy (anonymous):

i got 9y-2z=-30

OpenStudy (amistre64):

.........so did I but my fingers aint cooperating with me brain :)

OpenStudy (amistre64):

im gonna have to hold them under the iron to teach them a lesson....

OpenStudy (anonymous):

LOL, 8y+y would be 9y right?

OpenStudy (amistre64):

Yes, I agree........but that pinky finger is revolting ;)

OpenStudy (anonymous):

ok so next step is?

OpenStudy (amistre64):

well, since we already determined a solid value for "y"...lets use it ....wiat for it........NOW!!

OpenStudy (amistre64):

8(-2) -2z = -30 -16 - 2z = -30 -2z = -14 z = -14/-2 = 7 z = 7 yay!! write that down after you check me work..

OpenStudy (anonymous):

we subtract -16 from both sides

OpenStudy (amistre64):

why would we do that? -16-16 -2z = -30-16 is true, but it gets us nowhere fast. if we "add" 16 to both sides we get: -16+16 -2z = -30+16

OpenStudy (anonymous):

i got z=7

OpenStudy (amistre64):

me too :)

OpenStudy (amistre64):

y = -2, z = 7 lets use these in any of the others equations, but equation 1 seems lonely and wants some attention...so we will use it :) x + y + z = 4 x +-2 + 7 = 4 x + 5 = 4 x = 4-5 x = -1; y=-2; z=7 is the solution to our equations.

OpenStudy (amistre64):

you agree? we can double check by using these values in all the equations and seeing if the results are true...

OpenStudy (anonymous):

i agree

OpenStudy (amistre64):

its alot of mechanicl work, but once you know the underlying principles; the rest is just drudging thru the motions..

OpenStudy (anonymous):

something isnt right

OpenStudy (anonymous):

my school using mathlab and it says the answer is wrong x,y,z -1,-2,7

OpenStudy (anonymous):

it was 9y instead of 8y thats y i kept getting confused

OpenStudy (amistre64):

x+y+z=4 <- we know this ones good.... we used it to find x ...................................... 2(-1)+4(-2)+2(7)=4 -2 -8 +14 = 4 -10 + 14 = 4 4 = 4 ..... its good there ........................................ -(-1)+8(-2)-3(7)=-34 1 -16 -21 = -34 -14-21 = -34 -35 = -34 is off.....

OpenStudy (anonymous):

the equation is -x+8y-3z=-34 but when we added that to x+y+z=4 we got 0+8y-2z=-30 it should have been 0+9y-2z=-30

OpenStudy (amistre64):

1-16 = -15......

OpenStudy (amistre64):

you telling me you gave me the wrong equations?

OpenStudy (amistre64):

9y-2z=-30...and we used 8y right? lol

OpenStudy (amistre64):

this is soooo much easier to keep track of on paper, i gotta tell ya :)

OpenStudy (anonymous):

no the equations were right the math was wrong your fingers really screwed it up but just by one number the answer is x,y,z -1,-2,6

OpenStudy (anonymous):

add an extra 5mins on that iron

OpenStudy (amistre64):

lol.... gonna buy me one of those cigar cutters..... see what they think o that :)

OpenStudy (anonymous):

lmao, so what are you majoring in

OpenStudy (amistre64):

major in math hopefully; and minor in business

OpenStudy (anonymous):

no wonder i just thought you were a genius with math

OpenStudy (amistre64):

nah....just a normal genuis ;)

OpenStudy (amistre64):

gonna take calculus 1 and stats durig the summer if the classes dont cancel